29. Ethanol has an enthalpy of vaporization of 42.3 k.U/mol. The compound has a vapor pressure of 1.00 atm at 78.3 °C. At what temperature is the vapor pressure equal to 0.800 atm? (R 8.314 JK.mol) a. -83.8 °C b. -24.4 oC c. 62.6 C 73.0 OC e. 78.0 °C mmHg. What is the normal boiling point of this liquid? (R- 8.314 Jl(K mol)) d. 253 K. e. 234 K by 316 K. c. 269 K a. 287 K 31. Sulfur dioxide has a vapor pressure of 462.7 mm Hg at -21.0 oC and a vapor pressure of 140.5 mm Hg at -44.0 °C. What is the enthalpy of vaporization of sulfur dioxide? (R 8.314 J/K-mol) a. 0.398 kJ/mol b. 6.33 kJ/mol c. 14.0 kJ/mol Cd 24.9 /mol 32. Mount Everest rises to a height of 8.850 x 103 m above sea level. At this height, the atmospheric pressure is 231 mm Hg. At what temperature (in oC) does water boil at the summit of Mount Everest? JK mol) d.. 87 °C b, 69.0 °C C. 72 °C e, 364 °C a. 4.07 oC
Can you show the work for these problems?
Answer #1

Ans. 30. Vapor pressure for normal boiling = 1.00 atm = 760 mmHg

Putting the values in equation 1-

ln (760 mmHg/ 102 mmHg) = [30400 J mol-1 / 8.314 J K-1mol-1] [(1/269K) – (1/T2)]

Or,       2.00834562 = 3656.48 K [0.00372 K‑1 – (1/T2)]

Or,       (2.00834562 / 3656.48 K) = 0.00372 K‑1 – (1/T2)

Or,       0.0005493 K-1 – 0.00372 K-1 = – (1/T2)

Or,       T2 = 1/ (0.0031707 K-1) = 315.38 K

Hence, the normal boiling point (Temperature at which vapor pressure = 1.0 atm) = 315.38 K

            Note: Using 00C = 273.00 K may give a value much closer to 316 K

Ans. 31. Putting the values in equation 1

ln (462.7 mmHg/ 140.5 mmHg) = [dH / 8.314 J K-1mol-1] [(1/ 229.15K) – (1/ 252.15K)]

Or,       1.1918714071= [dH / 8.314 J K-1mol-1] x (0.004364 K‑1 – 0.003966 K-1)

Or,       1.1918714071= [dH / 8.314 J K-1mol-1] x (0.000398 K-1)

Or,       1.1918714071 / 0.000398 K-1 = dH / 8.314 J K-1mol-1

Or,       (2994.65 K-1) x 8.314 J K-1mol-1= dH

Or,       dH = 24897.53 J mol-1 = 24.9 kJ mol-1                                   ; [1 J = 10-3 kJ]

Hence, heat of vaporization = 24.9 kJ mol-1

Ans. 32. Putting the values in equation 1

ln (231 mmHg/ 760.0 mmHg) = [40700 J mol-1 / 8.314 J K-1mol-1] [(1/ 373 K) – (1/ T2)]

Or,       - 1.19= [4895.36 K-1] x (0.002681 K‑1 – 1/T2)

Or,       - 1.19/ 4895.36 K-1 =0.002681 K‑1 – 1/T2

Or,       - 0.000243 K-1 - 0.002681 K‑1 = – 1/T2

Or,       T2 = 1 / (0.002924 K-1) = 341.99 K                 

Hence, T2 = 341.99 K = 68.990C     

Note and Cautions:

1. Always write the values of T1 and P1 at their respective places. So, do it for T2 and P2.

However, it does not matter which set (of P1 and T1 ; or P2 and T2) of T and P you take as first and second conditions.

2. Temperature shall always be in K.

3. Note the sign. ln of a value less than 1.0 (say, 0.998) is a negative integer and that of greater than 1.0 is a positive integer.